A tea mixture was made from 40lbs of tea costing $5.00 per pound and 10lbs of tea costing $3.50 per pound. Find the cost of the tea mixture per pound.
Answer by checkley75(3666)

40*5.00+10*3.50=50X
200.00+35.00=50X
235.00=50X
X=235.00/50
X=$4.70 PRICE PER POUND FOR THE MIX.